OpenStudy (anonymous):

simplify and write fraction in lowest terms. 13/18 divided by 20/27 show work

OpenStudy (lollylau):

(13/18)*(27/20) =151/360

OpenStudy (anonymous):

13/18*27/20 =39/40

jhonyy9 (jhonyy9):

(13/18 )*(27/20)

jhonyy9 (jhonyy9):

13*3 ----- 2*20